OpenStudy (anonymous):

A recipe that makes 7 servings calls for 1 1/6 cups of juice. How many cups of juice are there per serving?

OpenStudy (studygurl14):

\[1\frac{1}{6} \div 7 = \frac{7}{6} \times \frac{1}{7} = answer\]

OpenStudy (anonymous):

7 - 42?

OpenStudy (studygurl14):

hint: the 7 in the numerator and the denomerator cancel each other out

OpenStudy (studygurl14):

Yes, that is correct.

OpenStudy (studygurl14):

But an easier way is to do this: